# LV-Eval: A Balanced Long-Context Benchmark with

### v1

@@ Abstract
State-of-the-art large language models (LLMs) are now claiming remarkable supported context lengths of $256k$ or even more. In contrast, the average context lengths of mainstream benchmarks are insufficient ($5k$-$21k$), and they suffer from potential knowledge leakage and inaccurate metrics, resulting in biased evaluation. This paper introduces LV-Eval, a challenging long-context benchmark with five length levels ($16k$, $32k$, $64k$, $128k$, and $256k$) reaching up to $256k$ words. LV-Eval features two main tasks, single-hop QA and multi-hop QA, comprising 11 bilingual datasets. The design of LV-Eval has incorporated three key techniques, namely confusing facts insertion, keyword and phrase replacement, and keyword-recall-based metric design. The advantages of LV-Eval include controllable evaluation across different context lengths, challenging test instances with confusing facts, mitigated knowledge leakage, and more objective evaluations. We evaluate 10 LLMs on LV-Eval and conduct ablation studies on the techniques used in LV-Eval construction. The results reveal that: (i) Commercial LLMs generally outperform open-source LLMs when evaluated within length levels shorter than their claimed context length. However, their overall performance is surpassed by open-source LLMs with longer context lengths. (ii) Extremely long-context LLMs, such as Yi-6B-200k, exhibit a relatively gentle degradation of performance, but their absolute performances may not necessarily be higher than those of LLMs with shorter context lengths. (iii) LLMs’ performances can significantly degrade in the presence of confusing information, especially in the pressure test of “needle in a haystack”. (iv) Issues related to knowledge leakage and inaccurate metrics introduce bias in evaluation, and these concerns are alleviated in LV-Eval. All datasets and evaluation codes are released at: https://github.com/infinigence/LVEval.
@@ 1 Introduction
Large language models (LLMs) have demonstrated exceptional performance on a variety of natural language processing tasks. The ability of long-context understanding is crucial for LLMs to deal with tasks based on longer contexts, such as books, lengthy chat history, and so on. Recently, extensive efforts have been devoted in enlarging the supported context length (i.e., the number of tokens that the model can accept as input) of LLMs. These efforts have pushed the supported context length of LLMs from $2k$ tokens to $32k$ tokens [1, 2, 3, 4, 5], and some models have achieved a remarkable context length of $128k$ and $200k$ [6, 7, 8].
In contrast to the rapid evolution of the models’ supported context length, existing benchmarks have lagged behind. The average word count in current long-context benchmarks typically falls within the range of $32k$ [9, 10, 11, 12, 13], considerably shorter in text length compared to the supported context lengths of state-of-the-art long context models. Moreover, previous benchmarks primarily consist of unaltered public documents and articles. This could be problematic for two reasons: (i) these data could potentially be involved in LLMs’ training processes, and (ii) the facts within them might be common-sense facts found in other training resources. The presence of this issue, known as “knowledge leakage” [14], can lead to models answering questions with memorization or common-sense knowledge instead of understanding long-range contexts. Last but not least, the automatic metrics employed in most of the existing benchmarks are susceptible to the variations in answer format and the inclusion of irrelevant words. Such metrics struggle to accurately assess the answer quality.
@@
ZeroSCROLLS [12] 10 13,556 1,023/320,774 none en
LooGLE [10] 7 21,247 10,927/246,182 none en
L-Eval [11] 20 12,993 2,119/170,256 none en
BAMBOO [13] 10 5,067 229/14,858 $4k$,$16k$ en+zh
LongBench [9] 21 9,486 128/71,954 00-$4k$,$4k$-$8k$,$8k$+ en+zh
@@
  • Keyword and phrase replacement to mitigate knowledge leakage: To mitigate the biased evaluation of long-context ability caused by knowledge leakage, we apply keyword and phrase replacement in the context and QA pairs. The replacement rules are annotated by human annotators. In this way, LV-Eval requires LLMs to rely on their understanding of the long context to answer questions rather than relying on memorization or common-sense knowledge.
@@
Utilizing LV-Eval, we evaluate 10 representative open-source and commercial LLMs with varying supported context lengths. Our findings are summarized as follows: (i) Commercial LLMs generally outperform open-source LLMs when evaluated within length levels shorter than their claimed context length. However, their overall performance is surpassed by open-source LLMs with longer context lengths. (ii) Extremely long-context LLMs, such as Yi-6B-200k, exhibit a relatively gentle degradation of performance, but their absolute performances may not necessarily be higher than those of LLMs with shorter context lengths. (iii) LLMs’ performances can significantly degrade in the presence of confusing information, especially in the pressure test of “needle in a haystack”. (iv) LLMs perform better on the single-hop QA task compared to the multi-hop QA task.
@@
Single-hop QA loogle-SD-mixup en 160 800
@@
lic-mixup zh 197 985
Table 2: Data statistics of LV-Eval. The abbreviations “CFI”, “KPR”, “AK” stand for “Confusing Fact Insertion”, “Keyword and Phrase Replacement”, and “Answer Keywords”, respectively. “#KPR” is the number of keyword and phrase replacement rules. Note that in factrecall-en and factrecall-zh, all QA pairs are the same across all test instances, i.e., there is only one unique QA pair for each of the two datasets.
@@ 2 Related Work
Considerable efforts have been devoted to enhancing the long-context processing capabilities of LLMs. One line of work focuses on making LLMs have extended context sizes without fine-tuning and behave normally on inputs longer than their training context lengths. The design and extrapolation method of the position encoding module [15, 16, 17] is crucial for this goal. Besides, several sparse attention techniques [18, 19] have also been proposed to avoid model collapse. These sparse attention techniques also alleviate the quadratic complexity w.r.t. the sequence length.
There are many other strategies aimed at enabling LLMs to effectively leverage long input contexts. The most commonly utilized strategy is long-context fine-tuning [20, 21, 22]. For instance, YaRN [22] conducts fine-tuning with $64k$ and $128k$ context lengths starting with Llama2-7B/13B, and Yi-6B-200k [8] is trained with $200k$ context length starting with its $4k$ variant. Other strategies include the recurrent- or memory-based architecture [23, 24, 25, 26, 27], and the retrieval- or summarization-based context compression techniques [28, 29, 9, 26], and so on.
@@
3 LV-Eval Benchmark
LV-Eval focuses on two types of QA tasks: single-hop QA and multi-hop QA, and is comprised of 11 QA datasets (6 in English and 5 in Chinese). The data statistics for LV-Eval are outlined in Table 2. Each test instance in LV-Eval comprises three parts: a context ($C$), a question ($Q$), and a GT answer ($A$). The context $C$ is a synthetic document containing the information required to answer the question $Q$.
Datasets in LV-Eval are constructed with existing public datasets as the source, except for factrecall-en and factrecall-zh, which are constructed using the data from PG19 [30] dataset and Dream of the Red Chamber book. Each dataset consists of five subsets of different lengths: $16k$, $32k$, $64k$, $128k$, and $256k$. All five subsets share the same question-answer (QA) pairs, meaning there are five contexts of varying lengths for each QA pair. This allows for a controllable evaluation of models’ long-context ability when testing the same set of questions with different context lengths. In total, LV-Eval comprises 1,729 QA pairs and 1,729$\times$5 = 8,645 synthetic contexts.
@@ 3 LV-Eval Benchmark
    Confusing facts insertion (Section 3.2.2): Then, in some datasets, we introduce confusing facts by generating them with GPT-4, manually revising them, and randomly inserting these into the context. These confusing facts bear similarities to the original supporting facts but are factually different, without contradicting the original information. This helps make the test instances more challenging.
@@
    Keyword and phrase replacement (Section 3.2.3): Finally, to reduce the impacts of knowledge leakage on evaluation results, we manually replace some keywords and phrases in the context and the QA pairs.
@@
Figure 2: Steps for confusing facts insertion. Firstly, we prompt GPT-4 to generate two descriptions that are close to the original fact. Then we ask human annotators to resolve any conflicts in the generated facts. For example, the first generated confusing fact “Albert Einstein was an Italian astronomer” is in conflict with the original fact and the human annotator revise it to “Albert Beverley was an Italian astronomer”. Finally, the confusing facts are inserted into a randomly position in the context.
@@ 3.1 Data Source and QA Pair Construction
We construct 11 datasets using public data sources, including Long-instruction-en2zh [31], HotpotQA [32], 2WikiMultihopQA [33], DuReader [34], LooGLE [10], LongBench [9], CMRC 2018 [35], MultiFieldQA [9], PG-19 [30] and the book of Dream of the Red Chamber). The construction of QA pairs in each of the 11 dataset is elaborated below.
Multi-hop QA.
In a multi-hop QA task, the reasoning to derive the answer needs to gather multiple pieces of information from various locations in the context. We construct five multi-hop QA datasets: dureader-mixup, loogle-CR-mixup, loogle-MR-mixup, hotpotwikiqa-mixup, and lic-mixup.
Single-hop QA.
In a single-hop QA task, only a single evidence in the context is needed to derive the answer. We construct six single-hop QA datasets: loogle-SD-mixup, cmrc-mixup, multifieldqa-en-mixup, multifieldqa-zh-mixup, factrecall-en, and factrecall-zh.
[…]
  • factrecall-en and factrecall-zh are two synthetic datasets designed to assess the LLMs’ ability to identify a small piece of evidence (“fact”) located at various locations within a lengthy context. More details are described in the appendix. As shown in Figure A10 A11, we write one English fact-question-answer pair for factrecall-en and one Chinese fact-question-answer pair for factrecall-ch. distracting documents are sourced from PG-19 dataset (English) and the book of Dream of the Red Chamber (Chinese) to create five contexts of different length levels. For each context, we generate 200 documents by inserting the fact at 200 evenly spaced positions within the context.
Figure 3: Steps for keyword and phrase replacement. First, given a QA pair, the annotators are asked to select keywords or phrases to replace and write a substitute for each. Then, the selected keywords and phrases are replaced throughout the entire original context and QA pair. Finally, annotators will check the modified context. If there is any conflict, the annotators are asked to revise the replacement rule until all conflicts are resolved.
@@ 3.2.1 Context Mixing Up
Can the LLMs identify the key evidences to answer the target question within a long context? To assess this ability, as shown in Figure 1, LV-Eval randomly mixes the supporting documents with various distracting documents to generate five contexts of varying length for a given QA pair. For 9 out of the 11 datasets (excluding factrecall-en and factrecall-zh), the distracting documents are chosen from the contexts corresponding to the non-selected QA pairs in the source dataset. For factrecall-en and factrecall-zh, the distracting documents are extracted from the PG-19 dataset and the book of Dream of the Red Chamber.
@@
Figure 4: Keyword-recall-based two-stage metric calculation. (a) The vanilla F1 score (red) is inflated high. With the keyword recall-based metric, the final score is set to zero due to the low recall of answer keywords. (b) The vanilla F1 score (red) is inflated high due to irrelevant words. By filtering of blacklisted words, the final score is better calibrated.
@@ 3.2.2 Confusing Facts Insertion
Can the LLMs identify the key evidences correctly if there are confusing facts in the context? To assess this ability, we apply confusing fact insertion in hotpotwikiqa-mixup, lic-mixup, multifieldqa-en-mixup, multifieldqa-zh-mixup, factrecall-en, and factrecall-zh, which inserts similar, factually different, non-contradictory facts into the context. These facts might mislead less meticulous models, leading them to generate incorrect answers.
The generation process of the confusing facts goes as follows. Firstly, we use the question and answer as the input, and prompt GPT-4 [7] to generate two descriptions that are close to the original fact. The prompt for GPT-4 is shown in Figure A12. Then, we ask human annotators to resolve any conflicts in the generated facts. As illustrated in Figure 2, the generated confusing fact “Albert Einstein was an Italian astronomer” is in conflict with the original fact. Therefore, the human annotator revise it to “Albert Beverley was an Italian astronomer”. After this generation and revising process, we insert the confusing facts into a randomly picked position between two sentences in the context.
@@ 3.2.3 Keyword and Phrase Replacement
Knowledge leakage is an important concern in LLM evaluation [14]. On the one hand, the test data are usually collected from open-access sources, and we cannot fully rule out the possibility of their being involved in some LLMs’ training process. On the other hand, some common-sense questions can be answered without referencing the provided context. Consequently, certain models might rely on memorization and common-sense knowledge to answer the questions rather than fully understanding the context. This will cause inflated benchmark scores to overrate the long-context ability of models.
To mitigate the influences of knowledge leakage on the evaluation results, we conduct keyword and phrase replacement according to manually crafted rules in hotpotwikiqa-mixup, cmrc-mixup, multifieldqa-en-mixup, multifieldqa-zh-mixup, factrecall-en, and factrecall-zh. Specifically, given a QA pair, the annotators are asked to select keywords or phrases for replacement and write a substitute for each. After the selected keywords and phrases are replaced throughout the entire context, the annotators review the modified context to check and resolve any conflicts: If there are conflicts, the annotators are asked to revise the replacement rule until all conflicts are resolved. One example of the keyword and phrase replacement process is shown in Figure 3. See Table 2 for the statistics of the number of replacement rules.
Model Name SFT Context Length HuggingFace / API Endpoint
Llama2-7B-Chat-hf [1] $4k$ meta-llama/Llama-2-7b-chat-hf
Qwen-7B-8k-Chat [3] $8k$ Qwen/Qwen-7B-Chat
Vicuna-7B-16k-v1.5 [2] $16k$ lmsys/vicuna-7b-v1.5-16k
ChatGLM3-6B-32k [4] $32k$ THUDM/chatglm3-6b-32k
Llama2-7B-32k-Instruct [36] $32k$ togethercomputer/Llama-2-7B-32K-Instruct
BlueLM-7B-32k-Chat [37] $32k$ vivo-ai/BlueLM-7B-Chat-32K
LongChat-7B-32k-v1.5 [36] $32k$ lmsys/longchat-7b-v1.5-32k
Yi-6B-200k [8] $200k$ 01-ai/Yi-6B-200K
GPT-4-8k [7] $8k$ gpt-4-0613
GPT-3.5-16k [38] $16k$ gpt-3.5-turbo-1106
Table 3: Information of evaluated LLMs.

### v2

@@ Abstract
State-of-the-art large language models (LLMs) are now claiming remarkable supported context lengths of $256k$ or even more. In contrast, the average context lengths of mainstream benchmarks are insufficient ($5k$-$21k$), and they suffer from potential knowledge leakage and inaccurate metrics, resulting in biased evaluation. This paper introduces LV-Eval, a challenging long-context benchmark with five length levels ($16k$, $32k$, $64k$, $128k$, and $256k$) reaching up to $256k$ words. LV-Eval features two main tasks, single-hop QA and multi-hop QA, comprising 11 bilingual datasets. The design of LV-Eval has incorporated three key techniques, namely confusing facts insertion (CFI), keyword and phrase replacement (KPR), and keyword-recall-based metric design. The advantages of LV-Eval include controllable evaluation across context lengths, challenging test instances with confusing facts, mitigated knowledge leakage, and more objective evaluation. We evaluate 15 LLMs on LV-Eval and conduct ablation studies on the benchmarking techniques. The results reveal that: (i) Moonshot-v1 and recent large-scale open-source models, such as Qwen-2.5-72B and Llama-3.1-70B, achieve the highest performance on LV-Eval, particularly at lengths below $64k$. (ii) Models exhibit distinct score trends. For example, GLM-4-9B-128k, Yi-6B-200k, and Llama3-8B-1M exhibit a relatively gentle degradation of performance, but their absolute performances may not necessarily be higher than those of LLMs with shorter context lengths. (iii) LLMs’ performances can significantly degrade in the presence of confusing information, especially in the pressure test of “needle in a haystack”. (iv) Issues related to knowledge leakage and inaccurate metrics introduce bias in evaluation, and these concerns are alleviated in LV-Eval. All datasets and evaluation codes are released at: https://github.com/infinigence/LVEval.
@@ 1 Introduction
Large language models (LLMs) have demonstrated exceptional performance on a variety of natural language processing tasks. The ability of long-context understanding is crucial for LLMs to deal with tasks based on longer contexts, such as books, lengthy chat history, and so on. Recently, extensive efforts have been devoted in enlarging the supported context length (i.e., the number of tokens that the model can accept as input) of LLMs. These efforts have pushed the supported context length of LLMs from $2k$ tokens to $32k$ tokens (Touvron et al., 2023; Zheng et al., 2023; Bai et al., 2023a; Zeng et al., 2022; Li et al., 2023a), and some models have achieved a remarkable context length of $128k$ and $200k$ (Peng et al., 2023a; Achiam et al., 2023; Yi, 2023).
In contrast to the rapid evolution of the models’ supported context length, existing benchmarks have lagged behind. The average word count in current long-context benchmarks typically falls within the range of $32k$ (Bai et al., 2023b; Li et al., 2023b; An et al., 2023; Shaham et al., 2023; Dong et al., 2023), considerably shorter compared to the supported context lengths of state-of-the-art long-context models. Moreover, previous benchmarks primarily consist of unaltered public documents and articles. This could be problematic for two reasons: (i) the data might be involved in LLMs’ training processes, and (ii) the facts within them might be common-sense facts found in other training resources. The presence of this issue, known as “knowledge leakage” (Zhou et al., 2023a), can lead to models answering questions with memorization or common-sense knowledge instead of understanding long-range contexts. Last but not least, the automatic metrics employed in most of the existing benchmarks are susceptible to the variations in answer format and the inclusion of irrelevant words. Such metrics struggle to accurately assess the answer quality.
@@
ZeroSCROLLS Shaham et al. (2023) 10 13,556 1,023/320,774 none en
LooGLE Li et al. (2023b) 7 21,247 10,927/246,182 none en
L-Eval An et al. (2023) 20 12,993 2,119/170,256 none en
BAMBOO Dong et al. (2023) 10 5,067 229/14,858 $4k$,$16k$ en+zh
LongBench Bai et al. (2023b) 21 9,486 128/71,954 00-$4k$,$4k$-$8k$,$8k$+ en+zh
@@
  • Keyword and phrase replacement to mitigate knowledge leakage: To mitigate the biased evaluation of long-context ability caused by knowledge leakage, we replace the keywords and phrases in the context and QA pairs. The replacement rules are annotated by human annotators. In this way, LV-Eval requires LLMs to rely on the understanding of context to answer questions rather than relying on memorization or common-sense knowledge.
@@
Findings.
We evaluate 15 LLMs on LV-Eval and summarize the main findings as follows: (i) Moonshot-v1 and recent large-scale open-source models, such as Qwen-2.5-72B and Llama-3.1-70B, achieve the highest performance on LV-Eval, particularly at lengths below $64k$. (ii) Models exhibit distinct score trends. For example, GLM-4-9B-128k, Yi-6B-200k, and Llama3-8B-1M exhibit a relatively gentle degradation of performance, but their absolute performances may not necessarily be higher than those of LLMs with shorter context lengths. (iii) LLMs’ performances can significantly degrade in the presence of confusing information, especially in the pressure test of “needle in a haystack”. (iv) Issues related to knowledge leakage and inaccurate metrics introduce bias in evaluation, and these concerns are alleviated in LV-Eval.
@@
Single-hop QA lic-mixup zh 197 985
loogle-SD-mixup en 160 800
@@
Table 2: Data statistics of LV-Eval. The abbreviations “CFI”, “KPR”, “AK” stand for “Confusing Fact Insertion”, “Keyword and Phrase Replacement”, and “Answer Keywords”, respectively. “#KPR” is the number of KPR rules. Note that in factrecall-en and factrecall-zh, all QA pairs are the same across all test instances, i.e., there is only one unique QA pair for each of the two datasets.
@@ 2 Related Work
The “needle in a haystack” (NIAH) task (Kamradt, 2023) has been very popular in assessing the long-context retrieval ability of LLMs. A recent benchmark, RULER (Hsieh et al., 2024), identifies the necessity of extending the NIAH task. It extends NIAH with diverse types and quantities of needles and distracting information. In our work, LV-Eval also includes an extended version of the NIAH task with our confusing fact insertion and keyword-and-phrase replacement techniques.
@@
Considerable efforts have been devoted to enhancing the long-context abilities of LLMs. One line of work focuses on making LLMs have extended context sizes without fine-tuning and behave normally on inputs longer than their training context lengths. The design and extrapolation method of the position encoding module (Su et al., 2024; Press et al., 2021; bloc97, 2023) is crucial for this goal. Besides, several sparse attention techniques (Han et al., 2023; Xiao et al., 2023) have also been proposed to avoid model collapse. These sparse attention techniques also alleviate the quadratic complexity w.r.t. the sequence length.
There are many other strategies aimed at enabling LLMs to effectively leverage long input contexts. The most commonly utilized strategy is long-context fine-tuning (Xiong et al., 2023; Li et al., 2023a; Peng et al., 2023b). For instance, YaRN (Peng et al., 2023b) conducts fine-tuning with $64k$ and $128k$ context lengths starting with Llama2-7B/13B, and Yi-6B-200k (Yi, 2023) is trained with $200k$ context length starting with its $4k$ variant. Other strategies include the recurrent- or memory-based architecture (Dai et al., 2019; Wu et al., 2021; Martins et al., 2022; Zhou et al., 2023b; Liang et al., 2023), and the retrieval- or summarization-based context compression techniques (Khandelwal et al., 2019; Borgeaud et al., 2022; Bai et al., 2023b; Zhou et al., 2023b), and so on.
@@ 3 LV-Eval Benchmark
LV-Eval focuses on two types of QA tasks: single-hop QA and multi-hop QA, and is comprised of 11 QA datasets (6 in English and 5 in Chinese). The data statistics for LV-Eval are outlined in Table 2. Each test instance in LV-Eval comprises three parts: a context ($C$), a question ($Q$), and a GT answer ($A$), where $C$ is a synthetic document containing the information required to answer $Q$.
Datasets in LV-Eval are constructed with existing public datasets as the source, except for factrecall-en and factrecall-zh, which are constructed using the data from PG19 (Rae et al., 2019) dataset and Journey to the West book. Each dataset consists of five subsets of different lengths: $16k$, $32k$, $64k$, $128k$, and $256k$. All five subsets share the same question-answer (QA) pairs, meaning there are five contexts of varying lengths for each QA pair. This allows for a controllable evaluation of models’ long-context ability when testing the same set of questions with different context lengths. In total, LV-Eval comprises 1,729 QA pairs and 1,729$\times$5 = 8,645 synthetic contexts.
@@
    Confusing Facts Insertion (CFI) (Section 3.2.2): Then, in some datasets, we introduce confusing facts by generating them with GPT-4, manually revising them, and randomly inserting these into the context. These confusing facts bear similarities to the original supporting facts but are factually different, without contradicting the original information. This helps make the test instances more challenging.
@@
    Keyword and Phrase Replacement (KPR) (Section 3.2.3): Finally, to reduce the impacts of knowledge leakage on evaluation results, we manually replace some keywords and phrases in the context and the QA pairs.
@@
During the benchmark construction, we employ and guide human annotators to revise the confusing facts, replace the keywords and phrases, and annotate the keywords in GT answers. The details of our annotation process are introduced in Appendix B.
@@ 3.1 Data Source and QA Pair Construction
We construct 11 datasets (see Table 2) using public data sources, including Long-instruction-en2zh (yuyijiong, 2023), HotpotQA (Yang et al., 2018), 2WikiMultihopQA (Ho et al., 2020), DuReader (Tang et al., 2020), LooGLE (Li et al., 2023b), LongBench (Bai et al., 2023b), CMRC 2018 (Cui et al., 2018), MultiFieldQA (Bai et al., 2023b), PG-19 (Rae et al., 2019) and the book of Journey to the West. The construction of QA pairs in each dataset is elaborated in Appendix A.
@@ 3.2 Context Construction
Figure 2: Steps for CFI. Firstly, we prompt GPT-4 to generate two descriptions that are close to the original fact. Then we ask human annotators to resolve any conflicts in the generated facts. For example, the first generated confusing fact “Albert Einstein was an Italian astronomer” is in conflict with the original fact and the human annotator revise it to “Albert Beverley was an Italian astronomer”. Finally, the confusing facts are inserted into a randomly position in the context.
@@ 3.2.1 Context Mixing Up
Can the LLMs identify the key evidences to answer the target question within a long context? To assess this ability, as shown in Figure 1, LV-Eval randomly mixes the supporting documents with various distracting documents to generate five contexts of varying length for a given QA pair. For 9 out of the 11 datasets (excluding factrecall-en and factrecall-zh), the distracting documents are chosen from the contexts corresponding to the non-selected QA pairs in the source dataset. For factrecall-en and factrecall-zh, the distracting documents are extracted from the PG-19 dataset and the book of Journey to the West.
@@ 3.2.2 Confusing Facts Insertion
Can the LLMs identify the key evidences correctly if there are confusing facts in the context? To assess this ability, we apply CFI in hotpotwikiqa-mixup, lic-mixup, multifieldqa-en-mixup, multifieldqa-zh-mixup, factrecall-en, and factrecall-zh, which inserts similar, factually different, non-contradictory facts into the context. These facts might mislead less meticulous models, leading them to generate incorrect answers.
The generation process of the confusing facts goes as follows. Firstly, we use the question and answer as the input, and prompt GPT-4 (Achiam et al., 2023) to generate two descriptions that are close to the original fact. The prompt for GPT-4 is shown in Figure A7. Then, we ask human annotators to resolve any conflicts in the generated facts. As illustrated in Figure 2, the generated confusing fact “Albert Einstein was an Italian astronomer” is in conflict with the original fact. Therefore, the human annotator revise it to “Albert Beverley was an Italian astronomer”. After this generation and revising process, we insert the confusing facts into a randomly picked position between two sentences in the context.
Figure 3: Steps for KPR. First, given a QA pair, the annotators are asked to select keywords or phrases to replace and write a substitute for each. Then, the selected keywords and phrases are replaced throughout the context and QA pair. Finally, annotators will check the modified context. If there is any conflict, the annotators are asked to revise the replacement rule until all conflicts are resolved.
@@ 3.2.3 Keyword and Phrase Replacement
Knowledge leakage is an important concern in LLM evaluation (Zhou et al., 2023a). On the one hand, the test data are usually collected from open-access sources, and we cannot fully rule out the possibility of their being involved in some LLMs’ training process. On the other hand, some common-sense questions can be answered without referencing the provided context. Consequently, LLMs might rely on memorization and common-sense knowledge to answer the questions rather than fully understanding the context. This will cause inflated benchmark scores to overrate the long-context ability of models.
To mitigate the influences of knowledge leakage on the evaluation results, we conduct KPR according to manually crafted rules in hotpotwikiqa-mixup, cmrc-mixup, multifieldqa-en-mixup, multifieldqa-zh-mixup, factrecall-en, and factrecall-zh. Specifically, given a QA pair, the annotators are asked to select keywords or phrases for replacement and write a substitute for each. After the selected keywords and phrases are replaced throughout the entire context, the annotators review the modified context to check and resolve any conflicts: If there are conflicts, the annotators are asked to revise the replacement rule until all conflicts are resolved. One example of the KPR process is shown in Figure 3. See Table 2 for the statistics of the number of replacement rules.